OpenStudy (anonymous):

Which of the following measurements is equal to 550000 µL?

OpenStudy (anonymous):

uL=10^-6 10^2 cL = 1 L 55 cL = 5.5 x 10^1 cL 5.5 x 10^1 cL x (1 L/10^2 cL) = 5.5 x 10^-1 L to check, I'll convert L to microL 5.5 x 10^-1 L x (10^6 microL/1 L) = 5.5 x 10^5 microL also written as: 550000 microL 55cL !!